im not sure why i cant pickout the most imporant info in these word problems please help...
If the length of a rectangle is 1km longer than twice the width then what is the length if the areaa is 55 square kilometer?

Answer by RAY100(1637)

Let L = Length,,,,,W= Width
.
Area = A = L*W,,,,,but L= 2W + 1
.
A= ( 2W+1) (W) = 2w^2 +W,,,,,and A= 55
.
2W^2 +W - 55 = 0,,,,,factor
.
(2W +11)(W-5) =0
.
2w+11=0,,,,W=-11/2,,,not realistic
.
W-5 =0,,,,W=5 and L= 2*5 +1 =11
.
.
check,,,,,,LW= 5*11=55,,,,,ok
.
2*5 +1 =11,,,,,,ok
.